OpenStudy (anonymous):

If f(x) = 2x - 3, g(x) = -2x^3 + 2 and h(x) = 3x, then h(2) + f(g(-1)) =

OpenStudy (anonymous):

please help

OpenStudy (a_clan):

11

OpenStudy (anonymous):

how u do?

OpenStudy (a_clan):

put x=2 in h(x)=3x to get h(2)

OpenStudy (a_clan):

then put x=-1 in g(x) to get g(-1). Subsequently, put the result of g(-1) in place of x in f(x) to get f(g(-1)). Then add both. h(2) + f(g(-1))

OpenStudy (anonymous):

so h(2)=6

OpenStudy (a_clan):

right

OpenStudy (anonymous):

f(g(-1)) =?

OpenStudy (anonymous):

2(-1)^3 -3 ?

OpenStudy (anonymous):

I see 2(-1)^3 +2 =5

OpenStudy (anonymous):

so 6+5=11

OpenStudy (anonymous):

no it is not right

OpenStudy (a_clan):

f(g(-1)) = f (result of g(-1)) = f (-2 *-1^3 +2) = f(2 + 2) = f(4) = 2(4)-3 = 5

OpenStudy (anonymous):

ah I understand now. Thank you

OpenStudy (a_clan):

:)